Suppose the coil of a meter has a resistance r of 20ohms and is fully deflected by a current of 5mA. Show that it can be converted to a meter of range 0-5A
1
Expert's answer
2019-08-29T10:11:47-0400

Let a shunt resistance of resistance "R_s" is connected in parallel to this meter,

Let "I" be the total current coming, "I_m" be the current passing through this meter and "I-I_m" be the current through shunt resistance

resistance of meter"=" 20"\Omega"

Voltage across shunt "=" "V_s=" Voltage across meter"=V_m"

Now "V_s=(I-I_m)\times R_s=I_m\times 20"

For caliberating this meter for given range "I" should be "5A", "I_m" should be "(5\times 10^{-3})A"

Now putting all values

"(5- 5\times10^{-3})\times R_s=(5\times 10^{-3})\times20"

"R_s=0.02\Omega"

Hence if a shunt of resistance "0.02\Omega" is connected in parellel to meter then meter will be caliberated to   range of 0-5"A"